OpenStudy (anonymous):

Can someone please help! What are the steps for solving 7x-17 ------- =4 for x? 8

OpenStudy (anonymous):

\[\frac{ 7x-17 }{ 8 }=4\]

OpenStudy (anonymous):

im guessing this is your equation

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok you want to get rid of the 8 first by using an operation that is opposite division.

OpenStudy (anonymous):

Ok, so I would multiply?

OpenStudy (anonymous):

Yes so you would multiply the eight on both sides of the equation so that the eights cancel on the left hand side where you x-variable is located.

OpenStudy (anonymous):

so what would you have now?

OpenStudy (anonymous):

56x-136 = 4?

OpenStudy (anonymous):

ok no you multiplied the eight on the left hand side and got rid ofthe lower 8 which is actually like multiplying by 64. you also did not multiply the 8 on the right hand side of the equals sign.

OpenStudy (anonymous):

\[(8)\frac{ 7x-17 }{ 8 } = 4(8)\] this is how it should look when you multiply the eights. the eight on the top will cancel with the eight on the bottom, and then you multiply 8 and 4 on the right side.

OpenStudy (anonymous):

\[7x-17=4(8) \] so then you are left with this.

OpenStudy (anonymous):

now see if you can finish it.

OpenStudy (anonymous):

Oh! ok! :D

OpenStudy (anonymous):

uhhh I got lost! :(

OpenStudy (anonymous):

where did you get lost write what you've done so far. we'll work it step by step.

OpenStudy (anonymous):

I'm not sure how to find x...

OpenStudy (anonymous):

ok well show me the work you've done after the equation I showed you.

OpenStudy (anonymous):

\[7x-17 =4(8)\] what did you do after this step?

OpenStudy (anonymous):

7x - 17 = 32

OpenStudy (anonymous):

ok good

OpenStudy (anonymous):

now remember the trick is to get x by itself.

OpenStudy (anonymous):

so what term do you think we want to move next?

OpenStudy (anonymous):

7?

OpenStudy (anonymous):

not yet. 7 is attached to 7 so we will leave those too together. Try again.

OpenStudy (anonymous):

im sorry attached to x

OpenStudy (anonymous):

17

OpenStudy (anonymous):

7 is attached to x so we will leave that term last.

OpenStudy (anonymous):

yes. that is the one

OpenStudy (anonymous):

but notice it is minus 17 so how do you think we can move -17 over to the right side of the equation?

OpenStudy (anonymous):

addition?

OpenStudy (anonymous):

exactly. so we will add 17 on both sides of the equation to cancel the 17 on the left.

OpenStudy (anonymous):

\[7x-17+17 = 32 +17\]

OpenStudy (anonymous):

so now what do you have?

OpenStudy (anonymous):

7x-49?

OpenStudy (anonymous):

oops!

OpenStudy (anonymous):

no not - 49 leave the 49 on the right side of the equals sign.

OpenStudy (anonymous):

7x=49

OpenStudy (anonymous):

yes exactly.

OpenStudy (anonymous):

so now what do you think will be the final step to solve for x?

OpenStudy (anonymous):

Divid 49 by 7?

OpenStudy (anonymous):

yes you got it.

OpenStudy (anonymous):

X=7

OpenStudy (anonymous):

so then what is the final answer?

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

yes that's it.

OpenStudy (anonymous):

well done

OpenStudy (anonymous):

Thank you sooo much! :D

OpenStudy (anonymous):

no problem.